# Fragile Preferences: A Deep Dive into Order Effects in Large Language Models

### v1

@@ Appendix A Color Comparison Sets and Prompts
Each model was asked to perform triplewise comparisons on four sets of three colors, categorized by tier. From best to worst, the tiers are Ideal, Fair, Plain, and Harsh. The method for generating these color sets is outlined in the Materials and Methods section of the main paper. Table 1 lists the specific colors used for each model and tier.
@@
The prompt used for triplewise comparisons is similar and omitted.
@@ Appendix B Resume Comparison Sets and Prompts
The selection methodology of resume for triple- and pairwise comparisons is outlined in the Materials and Methods section of the main paper. The prompts used to generate the resumes are the following.
prompts = [
@@ Appendix C Pairwise Color Comparisons
Table 2 reports the results at $T=1$, showing the proportion of times each position was chosen, along with effect sizes (Cohen’s $h$) and associated $p$-values. Effect sizes are measured against a 50% baseline, and $p$-values reflect the significance of deviation based on a two-sided binomial test. Even at the highest temperature setting used, most comparisons reveal statistically significant and substantial positional biases. We do not perform statistical tests for $T=0$ as the results are almost deterministic.
Figure 4 shows the results at different temperatures. At $T=0$ (the topmost plot), all models strongly prefer the first option for high-quality color sets, and shift to the second position for lower-quality sets. While higher temperatures introduce randomness, the positional bias remains robust across models and tiers.
Model Tier First / Second (%) Cohen’s $h$ $p$-value
@@
Table 2: Positional bias in pairwise color comparisons, $T=1$. Proportions of first and second selections across quality tiers. Cohen’s $h$ quantifies effect sizes from a 50% baseline. Asterisks indicate significance: ${\bf p<0.001}^{***}$.
@@ Appendix D Pairwise Resume Comparisons
For pairwise resume comparisons, we selected two resumes as outlined in Section 4 (Materials and Methods). We randomly selected four male names and four female names, and conducted exhaustive pairwise comparisons: $2$ genders $\times$ $\binom{4}{2}$ name pairs $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders, for a total of 48 comparisons per tier and profession.
@@
Figure 5 and Figure 6 plot the position effect results for the four professions, as well as aggregated results, at temperatures 0 and 1 respectively. The results align with findings from the color experiment: all models tend to prefer the first position for high-quality resumes but shift to favor the second position for low-quality ones. Figure 4 shows the corresponding results at $T=1$. Results for temperature 0.5 are consistent with the other findings and are omitted.
Model Tier First / Second (%) Cohen’s $h$ $p$-value
@@
Table 3: Pairwise position bias in resume comparisons, $T=1$. Results are aggregated across all professions, grouped by model and tier. For each condition, we report the proportion of times the first vs. second resume was chosen (First/Second %), the corresponding Cohen’s $h$ effect size, and the $p$-value from a two-sided binomial test against a 50% null. All p-values are from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ indicates statistical significance.
@@
Figure 5: Order effects in pairwise resume comparisons, $T=0$. Each panel shows the proportion of times each position was chosen for each profession (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated, broken down by LLM and quality tier.
@@
Figure 6: Order effects in pairwise resume comparisons, $T=1$. Each panel shows the proportion of times each position was chosen for each profession (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated, broken down by LLM and quality tier.
@@
As we reported in Appendix C, we conducted 50 comparisons for each permutation of each pair of colors at each temperature. Here, we conducted an additional 500 pairwise tests (250 iterations per permutation), for a total of 600 pairwise comparisons. The results are shown in Table 4. We found that in eight of the twelve tests, the model’s preference was statistically significant. Even when there is a clear preference, the effect size is small. This is to be expected, as the colors were chosen so that the LLM would be indifferent between them at $T=0$.
As we observed a strong name bias and moderately strong gender bias, we chose names in order to try to minimize the impact of these biases. We selected two female names that we found evoked a minimal bias in the pairwise comparisons: Ashley Williams and Emily Smith. Similarly to the color selection, we conducted 600 pairwise comparisons for resumes. Here, we needed to compare Resume 1 with Ashley Williams’ personal information to Resume 2 with Emily Smith’s personal information and vice versa. We therefore performed 150 iterations per permutation and resume-name pair, for a total of 600 pairwise comparisons. Here, the model’s preference was statistically significant in five of the twelve resume pairs. The results for resume comparisons are shown in Table 5.
Model Tier C1 / C2 (%) Cohen’s $h$ $p$-value
@@
Llama 3 8B Best 57 / 43 0.29 ${\bf 5.1\times 10^{-4}}^{***}$
@@
[…]
@@ Appendix F Triplewise Color Comparisons
We used a chi-square test to assess whether LLMs exhibit position bias when evaluating the three colors of equal quality. Under the null hypothesis of no positional preference, we would expect the model to select each position with equal probability (i.e., one-third each). A statistically significant deviation from this uniform distribution (as indicated by $p<0.001$) suggests that the model’s selection is systematically influenced by resume order rather than content. To quantify the strength of this effect, we report Cramér’s V as a measure of effect size. Table 6 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
Figure 7 shows the results at different temperatures. Similarly to the pairwise comparisons, while higher temperatures introduce randomness, the positional bias remains robust.
Model Tier First/Second/Third (%) Cramér’s $V$ $p$-value
@@
Table 6: Triplewise color comparisons, $T=1$. Proportions chosen from each position across models and quality tiers, reported as First/Second/Third (%). All $p$-values from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ denotes statistical significance.
@@ Appendix G Triplewise Resume Comparisons
The method for generating the resumes for triplewise comparisons is outlined in Section 4 (Materials and Methods). Similarly to the triplewise color comparisons, We conducted a chi-squared goodness-of-fit test under the null hypothesis that each position is selected with equal probability (i.e., 1/3 for each of the three positions), and measured effect sizes using Cramér’s V. Table 6 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
@@
Model Tier First / Second / Third (%) $\chi^{2}$ Cramér’s $V$ $p$-value
@@
Table 7: Order bias in triplewise resume comparisons, $T=1$. For each model and quality tier, “First/Second/Third (%)” shows the percentage of times the first, second, or third option was chosen. Position bias was evaluated with a chi-square goodness-of-fit test ($\chi^{2}$) against uniform choice; Cramér’s $V$ is the effect size. . Asterisks denote ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
@@
Figure 8: Positional Bias of triplewise resume comparisons, $T=0$. Positional order effects in triplewise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated across professions.
@@
Figure 9: Triplewise comparisons, $T=1$. Positional order effects in triplewise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated totals.
Appendix H Name Bias
H.1 Name Bias in Triplewise Comparisons
@@
Figure 10 visualizes the full distribution of name selection frequencies for each model using histograms with bin size 3. The expected number of times each name would be chosen if there was no name bias is 32. GPT-4o-mini chose each name between 26 and 38 times, consistent with little to no bias. Claude 3 Haiku ’s selections show high variance, with a wide spread and strong preference for certain names. Figure 11 shows box plots for each model, confirming these patterns: Claude 3 Haiku has the widest interquartile range and largest spread, while GPT-4o-mini’s selection counts are narrowly distributed with minimal outliers. Table 9 shows Claude 3 Haiku’s the top 5 least and most ‘favorite’ names: it chose some names as few as 8 times, and some as much as 62 times.
Model Gender $\chi^{2}$ $p$-value $n$
@@
H.2 Name Bias in Pairwise Comparisons
@@
Model Gender $\chi^{2}$ $p$-value Favorite Least Favorite
@@
Appendix I Gender Bias
We conducted cross‐gender pairwise comparisons using the same set of resumes and name–profile pairings employed in our same‐gender experiments. For each tier and profession, we compared every male–female pairing in both presentation orders, yielding 64 total comparisons per tier per profession (see Section 4). Table 11 reports the positional bias observed in these cross‐gender trials, while Table 12 reports the gender bias. Across nearly all models and tiers, positional effects exceed gender effects, with the sole exception of the ‘Best’ tier under Claude 3 Haiku.
To test whether this difference in effect sizes is itself statistically significant, we first defined, for each model–tier combination, the dominant position as the position (first or second) chosen most frequently in our same‐gender comparisons (Appendix D). We then aggregated two counts over the cross‐gender data: (1) the number of times the resume in that dominant position was selected, and (2) the number of times the female candidate was selected. The resulting positional and gender effect sizes are reported in Table 13. Finally, we performed a nonparametric bootstrap (10,000 resamples with replacement) on these aggregated counts to generate 95% confidence intervals for the difference in effect sizes; in every model–tier case the interval excluded zero, confirming that order effects are stronger than gender effects.
The reader may wish to compare Table 11 with Table 3, which gives the corresponding results for same-gender comparisons. On aggregate, GPT-4o-mini, Claude 3 Haiku and Llama 3 8B chose the candidate in the dominant position 93.6%, 87.6% and 89.2% of the time respectively, in same-gender comparisons, compared to 92.8%, 84.3% and 86.5% in cross-gender comparisons. We note that the name bias may play a role in these results; we do not attempt to disentangle the name and gender biases. Table 14 breaks down the results further, reporting the proportion of times the first position was chosen when the female/male candidate was presented first. Across all models and tiers, the first position was selected more often when the female candidate was presented first (except for one case, when the candidate in the first position was never selected).
@@ Appendix I Gender Bias
Model Tier First / Second (%) Cohen’s $h$ $p$-value
@@
Table 11: Pairwise position bias in cross-gender resume comparisons, $T=1$. Results are aggregated across all professions, grouped by model and tier. For each condition, we report the proportion of times the first vs. second resume was chosen (First/Second %), the corresponding Cohen’s $h$ effect size, and the $p$-value from a two-sided binomial test against a 50% null. All p-values are from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ indicates statistical significance.
Model Tier Female / Male (%) Cohen’s $h$ $p$-value
@@
[…]
@@
Model Tier Female when first (%) Male when first (%) Diff (%) $p$-value Cohen’s $h$
@@
Figure 12: Gender and position bias in pairwise comparisons, $T=0$. Gender effects in pairwise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated. Each panel shows the proportion of times each position was chosen, broken down by LLM and quality tier.
@@
Figure 13: Gender and position bias in pairwise comparisons, $T=1$. Gender effects in pairwise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated. Each panel shows the proportion of times each position was chosen, broken down by LLM and quality tier.

### v2

@@ Appendix A Color Comparison Sets and Prompts
Each model was asked to perform triplewise and pariwise comparisons on four sets of three colors, categorized by tier. From best to worst, the tiers are Ideal, Fair, Plain, and Harsh. The method for generating these color sets is outlined in the Materials and Methods section of the main paper. Table 1 lists the specific colors used for each model and tier.
@@ Appendix B Resume Comparison Sets and Prompts
The selection methodology of resume for triplewise and pairwise comparisons is outlined in the Materials and Methods section of the main paper. The prompt used to generate the resumes are the following.
prompt = [
@@ Appendix C Pairwise Color Comparisons
Table 2 reports the results at $T=1$, showing the proportion of times each position was chosen, along with effect sizes (Cohen’s $h$) and associated $p$-values. Effect sizes are measured against a 50% baseline, and $p$-values reflect the significance of deviation based on a two-sided binomial test. Even at the highest temperature setting used, most comparisons reveal statistically significant and substantial position biases. We do not perform statistical tests for $T=0$ as the results are almost deterministic.
[…]
@@
[…]
Figure 4 shows the results at different temperatures ($T=0,0.5,$ and $1$). At $T=0$ (the topmost plot), all models strongly prefer the first option for high-quality color sets, and shift to the second position for lower-quality sets. For instance, in color selection (4(a)), GPT-4o-mini displayed a primacy effect in the top two tiers (Ideal and Fair) and a recency effect in the bottom two (Plain and Harsh). Claude 3 Haiku and Llama 3 8B showed a primacy effect in the top one and three tiers, respectively, and a recency effect in the bottom three and one tiers, respectively.
While higher temperatures introduce randomness, the position bias remains robust across models and tiers.
@@
To assess whether the observed positional effects generalize beyond small-scale LLMs, we replicated our pairwise color comparison experiments using Claude 4 Sonnet, a frontier LLM with approximately 80 billion parameters—over 10× larger than Claude 3 Haiku. Across all four quality tiers, we observed consistent position biases similar to those reported for smaller models: primacy effects in higher-quality tiers and recency effects in lower-quality ones. These results indicate that position biases persist even in state-of-the-art models, supporting the claim that such effects are not limited to small-scale LLMs.
Figure 5: Order effects in pairwise comparisons at three temperature settings. The figure shows the percentage of selections by position (first or second), broken down by color tier for the Claude model across temperatures $T=0$, $0.5$, and 11 (left to right).
[…]
Ideal Robin’s Egg Blue Aqua 91 / 9 0.96 $3.3\times 10^{-18}$
Fair Gentle Coral Lilac 16 / 84 0.75 $2.6\times 10^{-12}$
Plain Ivory Eggshell White 4 / 96 1.17 $6.4\times 10^{-24}$
Harsh Onyx Obsidian 2 / 98 1.29 $1.6\times 10^{-26}$
Table 3: Pairwise position bias results for Claude 4 Sonnet (temperature = 1).
@@ Appendix D Pairwise Resume Comparisons
For pairwise resume comparisons, we selected two resumes as outlined in Section 4 (Materials and Methods). We randomly selected four male names and four female names, and conducted exhaustive pairwise comparisons: $2$ genders $\times$ $\binom{4}{2}$ name pairs $\times$ 22 name-resume assignments $\times$ 22 presentation orders, for a total of 48 comparisons per tier and profession.
@@
Figure 6 and Figure 7 plot the positional effect results for the four professions, as well as aggregated results, at temperatures 0 and 1 respectively. The results align with findings from the color experiment: all models tend to prefer the first position for high-quality resumes but shift to favor the second position for low-quality ones. Figure 7 shows the corresponding results at $T=1$. Results for temperature 0.5 are consistent with the other findings and are omitted.
Note that there is a difference between the pairwise color results at $T=0$ (4(a)) and the pairwise resume results (Figure 6). Specifically, in the color results, each color is chosen almost exactly 50% of the time. This matches the expectation that $\Pr[a\mid\{a,b\},0]$ will be very close to 0, $0.5$, or 11.666Recall that $\Pr[a\mid\{a,b\},t]$ is the overall probability of selecting $a$ when both permutations $(a,b)$ and $(b,a)$ are presented an equal number of times at temperature $t$.
At first glance, some pairwise resume comparisons seem to deviate from this pattern. The reason is that our resume dataset combines the same core content with different personal information. For example, let resumes X and Y be the core contents for the two resumes in the Good tier for Claude 3 Haiku, and let bl and cw be two male names. Xbl denotes resume X with the personal information of bl.
Because of the name bias, comparing Xbl with Ycw may yield different outcomes than comparing Xcw with Ybl. For example, if
$\Pr[\textsc{Xbl}\mid\{\textsc{Xbl},\textsc{Ycw}\},0]=1\quad\text{and}\quad\Pr[\textsc{Xcw}\mid\{\textsc{Xcw},\textsc{Ybl}\},0]=0.5,$
then core content X will be selected 75% of the time when results are aggregated over these two names. Since we use two core resumes and four names for each gender, this yields 12 $\{a,b\}$ pairs per gender, allowing a wide range of aggregate probabilities to emerge. Thus, deviations from exactly 0%, 50%, or 100% in the resume results are explained by the interaction between name bias and aggregation across name–content combinations. For more on the name bias, see Section I.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
As reported in Appendix C, we conducted 50 comparisons for each permutation of each pair of colors at each temperature. Here, we conducted an additional 500 pairwise tests (250 iterations per permutation), for a total of 600 pairwise comparisons. The results are shown in Table 5. In eight of the twelve tests, the model’s preference was statistically significant. Even when there is a clear preference, the effect size is small. This is to be expected, as the colors were chosen so that the LLM would be indifferent between them at $T=0$.
In order to try to minimize the impact of name and gender biases observed in pairwise resume comparisons, we selected two names that we found evoked minimal name biases when compared to each other: Ashley Williams and Emily Smith. Similarly to the color selection, we conducted 600 pairwise comparisons for resumes. Here, we needed to compare Resume 1 with Ashley Williams’ personal information to Resume 2 with Emily Smith’s personal information and vice versa. We therefore performed 150 iterations per permutation and resume-name pair, for a total of 600 pairwise comparisons. Here, the model’s preference was statistically significant in five of the twelve resume pairs. The results for resume comparisons are shown in Table 6.
[…]
@@
Table 5: Distinguishing tie-breaking from true distortions (colors). Values in the “C1/C2 (%)” column indicate the percentage of times each color was chosen the first color, where C1 and C2 denote the first and second color of that tier, denoted ‘Color 1’ and ‘Color 2’ in Table 1, respectively. Cohen’s $h$ quantifies effect size. Asterisks denote significance: $^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$.
@@
Llama 3 8B Best 57 / 43 0.28 $\bf 5.1\times 10^{-4}$$^{***}$
@@
Table 6: Distinguishing tie-breaking from true distortions (resumes). Values in the “R1/R2 (%)” column indicate the percentage of times each resume was chosen, where R1 and R2 denote the two resumes selected for pairwise comparisons. Cohen’s $h$ quantifies effect size. Asterisks denote significance: $^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$.
@@ Appendix F Triplewise Color Comparisons
We used a chi-square test to assess whether LLMs exhibit position bias when evaluating three colors of (approximately) equal quality. Under the null hypothesis of no positional preference, we would expect the model to select each position with equal probability (i.e., one-third each). Note that this is regardless of quality differentials: even if one color is significantly preferred to the others, we would expect each position to be selected one-third of the time. A statistically significant deviation from this uniform distribution (as indicated by $p<0.001$) suggests that the model’s selection is systematically influenced by resume order rather than content. To quantify the strength of this effect, we report Cramér’s V as a measure of effect size. Table 7 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
Figure 8 shows the results at different temperatures. Similarly to the pairwise comparisons, while higher temperatures introduce randomness, the position bias remains robust.
[…]
@@
[…]
@@ Appendix G Triplewise Resume Comparisons
The method for generating the resumes for triplewise comparisons is outlined in Section 4 (Materials and Methods). Similarly to the triplewise color comparisons, we conducted a chi-squared goodness-of-fit test under the null hypothesis that each position is selected with equal probability (i.e., one third for each of the three positions), and measured effect sizes using Cramér’s V. Table 7 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Appendix H Fourwise Color Comparisons
For fourwise comparisons, we manually curated a set of four colors, chosen to be roughly comparable, for each LLM. For each model, we presented all $4!=24$ permutations of these colors. We repeated each permutation four times at temperature 0, yielding 96 total queries per model.
The colors selected were:
  • GPT-4o-mini: Aqua Mist, Robin’s Egg Blue, Soft Sky Blue, Pale Turquoise,
  • Claude 3 Haiku: Pale Periwinkle, Pistachio, Blush Pink, Soft Terracotta,
  • Llama 3 8B: Emerald Green, Taupe, Navy Blue, Beige.
For GPT-4o-mini, the color set included its three colors from the Ideal tier, with Robin’s Egg Blue added. For Claude 3 Haiku, the colors came from its Fair tier, with Pistachio added. For Llama 3 8B, two colors were drawn from its Plain tier, with Emerald Green and Navy Blue added. Selection frequencies among the different colors were less balanced than in pairwise and triplewise comparisons, but each color was chosen at least once per repetition. GPT-4o-mini selected Aqua Mist and Robin’s Egg Blue 30 times each, Soft Sky Blue 24 times, and Pale Turquoise 12 times. Claude 3 Haiku most often selected Pale Periwinkle (42), followed by Pistachio (26), Blush Pink (24), and Soft Terracotta (4). Llama 3 8B favored Emerald Green (40), then Taupe (28), Navy Blue (20), and Beige (8).
Figure 11 shows the proportion of times each position (first, second, third, fourth) was chosen across all permutations, indicating consistent positional preferences in fourwise comparisons across all three models.
[…]
Appendix I Name Bias
I.1 Name Bias in Triplewise Comparisons
@@
Figure 12 visualizes the full distribution of name selection frequencies for each model using histograms with bin size 3. The expected number of times each name would be chosen if there was no name bias is 32. GPT-4o-mini chose each name between 26 and 38 times, consistent with little to no bias. Claude 3 Haiku ’s selections show high variance, with a wide spread and strong preference for certain names. Figure 13 shows box plots for each model, confirming these patterns: Claude 3 Haiku has the widest interquartile range and largest spread, while GPT-4o-mini’s selection counts are narrowly distributed with minimal outliers. Table 10 shows Claude 3 Haiku’s the top 5 least and most ‘favorite’ names: it chose some names as few as 8 times, and others as much as 62 times.
[…]
@@
I.2 Name Bias in Pairwise Comparisons
@@
[…]
@@
Appendix J Gender Bias
We conducted cross‐gender pairwise comparisons using the same set of resumes and name–profile pairings employed in our same‐gender experiments. For each tier and profession, we compared every male–female pairing in both presentation orders, yielding 64 total comparisons per tier per profession (see Section 4). Table 12 reports the position bias observed in these cross‐gender trials, while Table 13 reports the gender bias. Across nearly all models and tiers, positional effects exceed gender effects, with the sole exception of the Best tier under Claude 3 Haiku.
To test whether this difference in effect sizes is itself statistically significant, we first defined, for each model–tier combination, the dominant position as the position (first or second) chosen most frequently in the same‐gender comparisons (Appendix D). We then aggregated the following over the cross‐gender data: (1) the number of times the resume in that dominant position was selected, and (2) the number of times the female candidate was selected. The resulting positional and gender effect sizes are reported in Table 14. Finally, we performed a nonparametric bootstrap (10,000 resamples with replacement) on these aggregated counts to generate 95% confidence intervals for the difference in effect sizes; in every model–tier case the interval excluded zero, confirming that order effects are stronger than gender effects.
The reader may wish to compare Table 12 with Table 4, which gives the corresponding results for same-gender comparisons. On aggregate, GPT-4o-mini, Claude 3 Haiku and Llama 3 8B chose the candidate in the dominant position 93.6%, 87.6% and 89.2% of the time respectively, in same-gender comparisons, compared to 92.8%, 84.3% and 86.5% in cross-gender comparisons. We note that the name bias may play a role in these results; we do not attempt to disentangle the name and gender biases here. Table 15 breaks down the results further, reporting the proportion of times the first position was chosen when the female/male candidate was presented first. Across all models and tiers, the first position was selected more often when the female candidate was presented first (except for one case, when the candidate in the first position was never selected).  Figure 14 visually compares the results in  Table 12Table 4 and Table 15, by plotting the proportion of times the first resume was presented in 1) same-gender comparisons, 2) cross-gender comparisons, 3) cross-gender comparisons when the female candidate was presented first, and 4) cross-gender comparisons when the male candidate was presented first.
@@
[…]
@@
[…]
@@
Table 13: Gender bias at temperature 1. The two candidates differ in gender and alternate positions in the pairwise comparisons; values show the percentage of times the female vs. male candidate was selected, reported as Female/Male (%). Statistical significance is from binomial tests; effect sizes are Cohen’s $h$. Asterisks denote $^{*}p<.05$, $^{**}p<.01$, $^{***}p<.001$.
@@
Model Tier
Female when
first (%)
Male when
first (%)
[…]
@@
Figure 14: Same-gender vs cross-gender comparisons. Four results are displayed for comparison: 1) Order bias within same-gender comparisons, 2) Aggregate Order bias in cross-gender comparisons, 3) Order bias when the female candidate was the first position, and 4) Order bias when the male candidate was in the first option. In all cases, the plots report the proportion of times the first candidate presented was chosen.
@@
[…]
@@
[…]